### v2

@@ 4.1 One-factor Case
where we set $\phi=0.5$. We consider the dimensions $N=50,100,150,300$ and $500$, with sample sizes $T=200,500,800,1000$, and $1200$ for each $N$. Each loading value is generated independently from the uniform distribution $U(-2,2)$, and the $N$-dimensional loading vector is then re-normalized to have an $l_{2}$-norm strength of $\sqrt{N}$. For each configuration of $(N,T)$, we choose the sparsity $s=\left\lceil\sqrt{T}\right\rceil$, which is the nearest integer greater than or equal to $\sqrt{T}$, and we randomly generate $s$ integers from $\{1,...,T\}$ and keep those $s$ elements in $\underline{{\mathbf{f}}}_{1}=(f_{1},...,f_{T})^{\prime}$ to be nonzero, re-normalizing the sparse vector $\underline{{\mathbf{f}}}_{1}=(f_{1},...,f_{T})^{\prime}$ to have unit variance. We consider two scenarios for the idiosyncratic terms ${\mathbf{e}}_{t}$’s in each experiment:
@@
$d(\underline{\widehat{f}}_{1},\underline{f}_{1})=\sqrt{1-(\underline{\widehat{f}}_{1}^{\prime}\underline{f}_{1}/T)^{2}}.$ (26)
[…]
Table 1: The estimation accuracy of factors when the number of factors $r=1$. The results are based on the average of the measures defined in (26) through 500 replications. $N$ and $T$ denote the dimension and the sample size, respectively.
@@
$N$ 200 500 800 1000 1200 200 500 800 1000 1200
50 0.050 0.040 0.036 0.034 0.032 0.087 0.069 0.054 0.051 0.055
100 0.033 0.027 0.024 0.022 0.021 0.055 0.041 0.040 0.036 0.035
150 0.026 0.021 0.018 0.017 0.016 0.044 0.035 0.031 0.030 0.028
300 0.018 0.014 0.013 0.012 0.011 0.029 0.023 0.020 0.019 0.018
500 0.014 0.011 0.009 0.009 0.008 0.022 0.017 0.015 0.014 0.013
Next, we study the accuracy of the empirical recovery (ER) of the sparse factors. Let $S$ be the set of true indexes of the $s$ nonzero elements in $\underline{f}_{1}$, and $\widehat{S}$ the set of indices of the $\widehat{s}$ nonzero elements in $\underline{\widehat{f}}_{1}$. For each replication, define the empirical recovery rate of the non-sparse indices as
@@
where #$\#\{S\cap\widehat{S}\}$ is the cardinality of the intersection between the estimated indices and the true ones. Table 2 presents the empirical recovery rate of the sparsity in the factor process using the measure defined in (27). From Table 2 we see that the empirical accuracy increases as the the sample size increases for each fixed dimension $N$, and it also increases as the dimension $N$ increases for each fixed sample size $T$. This is also in agreement with our asymptotic results in Section 3.
Table 2: Empirical recovery rate of the sparsity in the factor process when the number of factors $r=1$. The empirical recovery rate in each replication is calculated as in (27). $N$ and $T$ denote the dimension and the sample size, respectively. 500 replications are used in the experiments.
@@
$N$ 200 500 800 1000 1200 200 500 800 1000 1200
50 0.918 0.933 0.936 0.938 0.940 0.884 0.900 0.914 0.916 0.910
100 0.940 0.950 0.954 0.956 0.957 0.916 0.930 0.930 0.935 0.937
150 0.952 0.959 0.962 0.963 0.965 0.931 0.939 0.943 0.945 0.947
300 0.969 0.971 0.971 0.974 0.973 0.951 0.956 0.958 0.962 0.961
500 0.971 0.978 0.979 0.980 0.980 0.957 0.965 0.969 0.970 0.971
Finally, we evaluate the distribution of the estimated loadings as described in Theorem 3. For simplicity, we consider the case when the noises are $i.i.d.$ and plot the empirical histogram of the estimated loading associated with the first series in Figure 1. The variance of the normal curve is based on the average of the estimated variance across 2000 replications. From Figure 1, we see that the estimators behave closely to normal, which is in agreement with our asymptotic theory.
Figure 1: Histograms of the estimated loading of $\sqrt{T}(\widehat{\boldsymbol{\lambda}}_{1}-{\mathbf{H}}_{s}\boldsymbol{\lambda}_{1})$, where we set $(N,T)=(300,500)$ and $s=\left\lceil\sqrt{T}\right\rceil$. The settings of the parameters are the same as those in Table 1. The curve represents the density of normal distribution where the standard errors are estimated by the sample versions using the residuals. 2000 replications are used in the experiments.
@@ 4.2 Multi-factor Case
where $\Psi=\mbox{diag}(0.5,-0.6,0.7)$ is a diagonal matrix. We consider the dimensions $N=50,100,150,200$, and $300$, with the sample size $T=100,200,300,500$, and $800$ for each $N$ in this experiment. For each configuration of $(N,T)$, we set the sparsity parameters to $s_{1}=s_{2}=s_{3}=\left\lceil\sqrt{T}\right\rceil$. We first randomly generate $s_{1}$ indexes from $\{1,...,T\}$ to form an index set $S_{1}$, such that we only keep the corresponding $s_{1}$ time points of $\underline{{\mathbf{f}}}_{1}$ and the remaining ones are set to zero. The index set $S_{2}$ is formed by generating $s_{2}$ indexes $\{1,...,T\}\setminus S_{1}$, keeping only those $s_{2}$ locations of $\underline{{\mathbf{f}}}_{2}$. the sparse $\underline{{\mathbf{f}}}_{3}$ is obtained by repeating the above procedure. Then each $\underline{{\mathbf{f}}}_{j}$ is normalized to have unit variance. For the generation of the loading matrix $\boldsymbol{\Lambda}\in R^{N\times r}$, we first generate an $N\times r$ matrix ${\mathbf{M}}$ with elements independently generated from $U(-2,2)$, we then perform a singular-value decomposition on ${\mathbf{M}}$, and the left singular matrix is ${\mathbf{U}}$. The loading $\boldsymbol{\Lambda}$ is taken as ${\mathbf{U}}$ multiplied by $\sqrt{N}\mbox{diag}(3,2,1)$ on its right. We also consider the two scenarios for the idiosyncratic terms as in Section 4.1. A total of 500 replications are used throughout the experiments.
@@
The average estimation errors of the factors over 500 replications are reported in Table 3. From Table 3, we observe a similar pattern to that in Section 4.1. For each fixed dimension $N$, the estimation error generally decreases as the sample size $T$ increases. Additionally, the error decreases as the dimension $N$ increases for each fixed $T$, which is in agreement with our asymptotic theory in Theorem 4, regardless of whether the idiosyncratic terms are i.i.d. or dynamically dependent.

### v3

@@ 4.1 One-factor Case
where we set $\phi=0.5$. We consider the dimensions $N=50,100,150,300$ and $500$, with sample sizes $T=200,500,800,1000$, and $1200$ for each $N$. Each loading value is generated independently from the uniform distribution U(2,2)U(-2,2), and the $N$-dimensional loading vector is then re-normalized to have an $l_{2}$-norm strength of $\sqrt{N}$. For each configuration of $(N,T)$, we choose the sparsity $s=T/10$, and we randomly generate $s$ integers from $\{1,...,T\}$ and keep those $s$ elements in $\underline{{\mathbf{f}}}_{1}=(f_{1},...,f_{T})^{\prime}$ to be nonzero, re-normalizing the sparse vector $\underline{{\mathbf{f}}}_{1}=(f_{1},...,f_{T})^{\prime}$ to have unit variance. We consider two scenarios for the idiosyncratic terms ${\mathbf{e}}_{t}$’s in each experiment:
@@
$d(\underline{\widehat{\mathbf{f}}}_{1},\underline{{\mathbf{f}}}_{1})=\sqrt{1-(\underline{\widehat{\mathbf{f}}}_{1}^{\prime}\underline{{\mathbf{f}}}_{1}/T)^{2}}.$ (25)
Table 1 reports the average estimation errors over 500 replications under two scenarios for the idiosyncratic components. Proposal refers to our proposed method, while Lasso denotes the Lasso-regularized approach applied to the factor process, following a similar procedure to Algorithm 1 in Kristensen (2017), with the tuning parameter set to $\psi_{T}=0.6f_{0}$, where $f_{0}$ is the minimum nonzero component (in absolute value) of $\underline{\bf f}_{1}$. As shown in Table 1, for each scenario, the estimation error decreases as the dimension $N$ increases for a fixed sample size $T$. In contrast, for fixed $N$, the error remains relatively stable or slightly increases as $T$ grows, which is consistent with the theoretical rate $\sqrt{\log(T)/N}$—a term that dominates $1/T$ for large TT—as established in Theorem 1. These results are consistent with our asymptotic theory. Moreover, the results clearly indicate that the Lasso procedure tends to incur additional estimation errors across all configurations of $(N,T)$. This highlights the advantage of the proposed hard-thresholding technique.
Table 1: The estimation accuracy of factors when the number of factors $r=1$. The results are based on the average of the measures defined in (25) through 500 replications. NN and $T$ denote the dimension and the sample size, respectively. The Proposal refers to our proposed method, while Lasso denotes the Lasso-regularized approach applied to the factor process, which is similar to Algorithm 1 in Kristensen (2017).
@@
Method NN 200 500 800 1000 1200 200 500 800 1000 1200
50 0.060 0.062 0.062 0.062 0.062 0.095 0.102 0.097 0.097 0.096
100 0.040 0.041 0.041 0.041 0.041 0.064 0.068 0.068 0.071 0.065
Proposal 150 0.031 0.033 0.033 0.033 0.033 0.053 0.054 0.055 0.054 0.054
300 0.021 0.022 0.022 0.022 0.022 0.035 0.036 0.035 0.035 0.035
500 0.016 0.016 0.012 0.017 0.017 0.026 0.027 0.027 0.027 0.027
50 0.264 0.200 0.185 0.176 0.172 0.312 0.261 0.240 0.225 0.229
100 0.245 0.176 0.152 0.144 0.137 0.273 0.214 0.193 0.193 0.177
Lasso 150 0.230 0.163 0.139 0.131 0.123 0.256 0.192 0.174 0.164 0.157
300 0.229 0.156 0.123 0.116 0.108 0.242 0.172 0.141 0.135 0.127
500 0.218 0.144 0.116 0.111 0.101 0.226 0.155 0.129 0.126 0.116
Next, we study the accuracy of the empirical recovery (ER) of the sparse factors. Let $S$ be the set of true indexes of the $s$ nonzero elements in $\underline{{\mathbf{f}}}_{1}$, and $\widehat{S}$ the set of indices of the $\widehat{s}$ nonzero elements in $\underline{\widehat{\mathbf{f}}}_{1}$. For each replication, define the empirical recovery rate of the non-sparse indices as
@@
where $\{S\cap\widehat{S}\}$ is the cardinality of the intersection between the estimated indices and the true ones. Table 2 presents the empirical recovery rate of the sparsity in the factor process using the measure defined in (26). Similar to the comparison in Table 1, we also compare the proposed method with the Lasso approach. From Table 2, we observe that empirical accuracy improves as the dimension $N$ increases for each fixed sample size $T$, which is consistent with the asymptotic results presented in Section 3. In addition, the Lasso procedure performs comparably to the proposed method in recovering the sparsity structure of the factor process, with the proposed method only slightly outperforming Lasso in certain cases. Taken together with the results in Table 1, this suggests that the hard-thresholding technique provides an improvement in accuracy while remaining easy to implement.
Table 2: Empirical recovery rate of the sparsity in the factor process when the number of factors $r=1$. The empirical recovery rate in each replication is calculated as in (26). NN and $T$ denote the dimension and the sample size, respectively. 500 replications are used in the experiments. The Proposal refers to our proposed method, while Lasso denotes the Lasso-regularized approach applied to the factor process, which is similar to Algorithm 1 in Kristensen (2017).
@@
Method NN 200 500 800 1000 1200 200 500 800 1000 1200
50 0.916 0.909 0.909 0.909 0.909 0.880 0.868 0.873 0.880 0.874
100 0.937 0.934 0.933 0.934 0.934 0.908 0.902 0.904 0.901 0.906
Proposal 150 0.947 0.944 0.944 0.944 0.944 0.919 0.919 0.918 0.918 0.917
300 0.962 0.961 0.959 0.959 0.958 0.943 0.941 0.941 0.940 0.941
500 0.969 0.969 0.968 0.967 0.968 0.956 0.952 0.952 0.955 0.953
50 0.916 0.909 0.909 0.909 0.909 0.878 0.867 0.872 0.879 0.873
100 0.937 0.934 0.933 0.934 0.934 0.908 0.901 0.903 0.901 0.906
Lasso 150 0.947 0.944 0.944 0.944 0.944 0.918 0.918 0.918 0.917 0.917
300 0.962 0.961 0.959 0.959 0.958 0.943 0.941 0.940 0.940 0.941
500 0.969 0.969 0.968 0.967 0.968 0.955 0.952 0.952 0.955 0.953
Finally, we evaluate the distribution of the estimated loadings as described in Theorem 3. For simplicity, we consider the case when the noises are $i.i.d.$ and plot the empirical histogram of the estimated loading associated with the first series in Figure S.2 of the Supplement. The variance of the normal curve is based on the average of the estimated variance across 2000 replications. From Figure S.2, we see that the estimators behave closely to normal, which is in agreement with our asymptotic theory.
@@ 4.2 Multi-factor Case
where $\Psi=\mbox{diag}(0.5,-0.6,0.7)$ is a diagonal matrix. We consider the dimensions $N=50,100,150,200$, and $300$, with the sample size $T=100,200,300,500$, and $800$ for each $N$ in this experiment. For each configuration of $(N,T)$, we set the sparsity parameters to $s_{1}=s_{2}=s_{3}=T/10$. We first randomly generate $s_{1}$ indexes from $\{1,...,T\}$ to form an index set $S_{1}$, such that we only keep the corresponding $s_{1}$ time points of $\underline{{\mathbf{f}}}_{1}$ and the remaining ones are set to zero. The index set $S_{2}$ is formed by generating $s_{2}$ indexes $\{1,...,T\}\setminus S_{1}$, keeping only those $s_{2}$ locations of $\underline{{\mathbf{f}}}_{2}$. the sparse $\underline{{\mathbf{f}}}_{3}$ is obtained by repeating the above procedure. Then each $\underline{{\mathbf{f}}}_{j}$ is normalized to have unit variance. For the generation of the loading matrix $\boldsymbol{\Lambda}\in R^{N\times r}$, we first generate an $N\times r$ matrix $\mathbf{M}$} with elements independently generated from U(2,2)U(-2,2), we then perform a singular-value decomposition on $\mathbf{M}$}, and the left singular matrix is $\mathbf{U}$}. The loading $\boldsymbol{\Lambda}$ is taken as $\mathbf{U}$} multiplied by $\sqrt{N}\mbox{diag}(3,2,1)$ on its right. We also consider the two scenarios for the idiosyncratic terms as in Section 4.1. A total of 500 replications are used throughout the experiments.
@@
The average estimation errors of the factors over 500 replications are reported in Table S.I of the Supplement. From Table S.I, we observe a similar pattern to that in Section 4.1. For each fixed sample size $T$, the error decreases as the dimension $N$ increases, which is in agreement with our asymptotic theory in Theorem 4, regardless of whether the idiosyncratic terms are i.i.d. or dynamically dependent.
where $s=T/10$ and $\widehat{S}_{j}$ is the the estimated nonzero locations in $\underline{\widehat{\mathbf{f}}}_{j}$ for $j=1,2$, and $3$. Table S.II of the Supplement reports the empirical accuracy of the sparsity locations in the factor processes when $r=3$. From Table S.II, we see that the pattern is also similar to the case when $r=1$ in Section 4.1. The empirical results are in line with our asymptotic ones in the sense that the estimation accuracy will increase as the dimension increases for each fixed $T$.
Although the ratio-based method of (12) for determining the number of factors has been shown to be valid in many previous studies, we conduct an auxiliary experiment to verify its efficacy under our setting. Table S.III of the Supplement reports the empirical probabilities of $P(\widehat{r}=r)$ in 500 replications under the aforementioned setting. From Table S.III, we see that the ratio-based method performs well, with most of the empirical probabilities being close to one. This is understandable since all the factors are strong ones. Similar results can be found in Gao and Tsay (2023).
4.3 Determining the Sparsity with Cross-Validation
In this section, we study the estimation accuracy of the information criterion discussed in Section 2.4 and Remark 4 in estimating the sparsity of the factors. For simplicity, we only consider the case when $r=1$, but similar results can be obtained for $r>1$. The generation of the factors and the data is similar to those in Section 4.1, but we only keep the largest $s=T/10$ elements of the factor process in absolute value for each sample size $T$. We consider the dimension $N=50,150,200$, and $300$, and the sample size $T=100,200,300,500$, and $800$ for each $N$ in this section. For each configuration of $(N,T)$, we set the number of partitions $J=1$ for simplicity and set $N_{1}=N_{2}=N/2$ in the partition. Since $s=T/10$ is diverging with the sample size $T$, we will use the information criterion defined in (24) to estimate the sparsity. 500 replications are used throughout the experiment. Table S.IV of the Supplement reports the Empirical probabilities (EP) of determining the sparsity parameter using the information criterion in (24) when the number of factors $r=1$. The empirical probabilities are calculated based on the 500 experiments. From Table S.IV we see that that the proposed information criterion estimates the sparsity parameter effectively. While dynamic dependence in idiosyncratic terms reduces accuracy, its performance generally improves with increasing $N$ for each fixed $T$, which is consistent with Theorem 4.
5 An Empirical Application to Stock Returns
5.1 Data
Similar to Pelger (2020), we group these 332 stocks into 14 categories by industry sector, as shown in Table S.V of the Supplement. The data encompasses a wide range of industry sectors within the stock market, including Oil, Finance, Electricity, Technology, Food, Manufacturing, Pharma & Chemicals, Primary Manufacturing, Machinery, Health, Transportation, Trade, Services, and Mining.
5.2 Sparse Factor Estimation
We first determine the number of factors for the centered data $\mathbf{X}$} using the eigenvalue-ratio method in (12). Figure S.3 of the Supplement plots the ratios of eigenvalues of ${\mathbf{X}}{\mathbf{X}}^{\prime}$ and shows that the largest gap between the eigenvalues occurs between $\widehat{\lambda}_{1}$ and $\widehat{\lambda}_{2}$, implying that the number of factors is $\widehat{r}=1$. Next, we apply Algorithm 1 and the cross-validation method in Section 2.4 with the number of random partitions $J=10$, where the seed number is set to be 1234 in R. For the sparsity of the factors over the time horizon, we conduct a grid search over the time interval $s\in[\lceil\frac{T}{18}\rceil-50,\lceil\frac{T}{12}\rceil+50]=[132,232]$, where $\lceil x\rceil$ denotes the smallest integer that exceeds $x$. Figure 2 plots the information criterion $IC(s)$ defined in (24) and the estimated $\widehat{s}=211$. This suggests that there are 211 days during which stock returns were more strongly influenced by significant systematic risks. We also searched over the neighborhoods of $T/10$ and $T/6$, and found that the minimum of $IC(s)$ in these regions is greater than $IC(211)$. Moreover, the entire path of $IC(s)$ increases after $s=211$, indicating that $\widehat{s}=211$ is a suitable estimate. It is important to note that this does not imply the absence of systematic risk on the remaining days; rather, the method is designed to highlight periods with stronger and more dominant risk signals. Compared to the full time span of $T=3273$, this represents a substantial reduction in temporal dimensionality, thereby simplifying subsequent analysis.
Figure 2: Plots of the information criterion $IC(s)$, defined in equation (24) of the main text, over the range $s\in[\lceil\frac{T}{18}\rceil-50,\lceil\frac{T}{12}\rceil+50]=[132,232]$. We use $J=10$ cross-validations for each $s$.
We then plot the estimated sparse factor returns in Figure 3. Similar to the example in the Introduction, we find that ${\sum_{t=1}^{3273}(f_{t}^{s})^{2}}/{\sum_{t=1}^{3273}f_{t}^{2}}=62.5\%$, indicating that the 211 nonzero factor values—representing only 6.4% of the total 3273 time points—account for 62.5% of the variance in the non-sparse factor process ${\mathbf{f}}_{t}$. Moreover, the plot reveals approximately three distinct clusters of time periods associated with significant systematic risks affecting the stock market. The first cluster is centered around 2008, driven by the 2007–2008 financial crisis. The second and third clusters are closely related, both associated with the European sovereign debt crisis. The most prominent cluster is centered around 2008, driven by the 2007–2008 financial crisis. This period saw severe disruptions in the financial markets, with significant declines in stock returns across the board. For the two closely related clusters associated with the European sovereign debt crisis, the first cluster corresponds to the initial phase of the crisis in 2010, while the second cluster relates to the intensification of the crisis in 2011 and 2012.
Next, we estimate the loadings of the 332 stock returns given the sparse common factors. We present the estimated loadings, indicating the level of dependence of 14 different industry sectors on the common risk factor in Figure 4. The loadings quantify the extent to which each sector is influenced by these common factors. Here are the detailed observations based on the average dependence values: Oil Sector: Exhibits moderate dependence on common risk factors, with an average loading of 0.0125. Finance Sector: Shows significant dependence, with an average loading of 0.0165. Electricity Sector: Demonstrates relatively low dependence, with an average loading of 0.0062. Technology Sector: Reflects moderate dependence, with an average loading of 0.0097. Food Sector: Indicates low dependence on common factors, with an average loading of 0.0046. Manufacturing Sector: Displays moderate dependence, with an average loading of 0.0116. Pharma & Chemicals Sector: Shows relatively low dependence compared to other sectors, with an average loading of 0.0070. Primary Manufacturing Sector: Exhibits moderate dependence, with an average loading of 0.0133. Machinery Sector: Demonstrates moderate dependence, with an average loading of 0.0109. Health Sector: Indicates relatively low dependence, with an average loading of 0.0061. Transportation Sector: Shows moderate dependence, with an average loading of 0.0099. Trade Sector: Reflects moderate dependence, with an average loading of 0.0084. Services Sector: Exhibits moderate dependence, with an average loading of 0.0111. Mining Sector: Indicates relatively low dependence on common risk factors, with an average loading of 0.0079.
5.3 Bridging Time and Risk Factors
To further identify the systematic factors over time, referred to as time factors, Table S.VII in Sec. D of the Supplement provides a comprehensive list of dates with significant systematic risk factors, the reasons for these factors, and their associated time factors. Each entry explains the specific event and its impact on stock returns, summarized according to the descriptions of the time factors in Table S.VI of the Supplement. These reasons are extracted from the daily reports on CNN Money (www.money.cnn.com) after the market closes on each trading day. This website was shut down in late 2018 and its content was merged into the main CNN Business section.
We also plot the frequency charts of the nine factors over the time horizon in Figure 5. The most frequently mentioned factor is market sentiment, highlighting its dominant role in driving stock price fluctuations. Economic indicators and government policies follow closely, indicating the significant impact of macroeconomic data and policy decisions on market behavior. Company-specific factors also play an important role, affecting individual stock performance and, by extension, broader market trends. Factors related to Europe, such as economic conditions and crises, have a moderate influence, reflecting the interconnectedness of global markets. Oil price fluctuations, China’s economic activities, and global events are also notable contributors, underscoring the importance of global economic dynamics. Credit risk, while mentioned less frequently, remains a critical factor during periods of financial instability. Further implications of the results are discussed in Section B.1 of the Supplement for brevity.
Figure 5: Frequency of Time Factors Influencing Stock Market Movements from January 2004 to December 2016. The chart visualizes the occurrence of various time factors, including market sentiment, economic indicators, government policies, company-specific factors, Europe, oil price, China, global events, and credit risk, based on their frequency of mention in Table S.VII of the Supplement.
6 Conclusion
Sparse factor modeling aims to extract meaningful insights by identifying a subset of relevant variables, enhancing interpretability while reducing computational complexity. This study presents a novel perspective by assuming that factor processes exhibit sparsity over time rather than imposing sparsity on loadings. This approach is particularly suited to scenarios where systematic co-movements are significant only during specific periods, such as financial crises or policy changes.
This paper introduces approximate factor models and a novel sparse asymptotic PCA (APCA) method, extending the asymptotic PCA framework of Connor and Korajczyk (1986, 1988) to high-dimensional settings. Theoretical properties of the proposed method, including the cross-sectional cross-validation approach for estimating sparsity, are rigorously established. Through simulation and empirical studies, the method is shown to be effective in delivering interpretable results and linking co-movements to specific events. This work not only advances the statistical methodology of sparse factor modeling but also establishes a meaningful connection between time-specific events and risk factors in economic and financial systems.
Supplementary Material
This supplementary material includes some examples of sparse factors over time horizons in finance and economics, additional tables and figures used in the numerical analyses, theoretical proofs of the main theorems, and a description of the text data extracted from CNN Money (www.money.cnn.com).
Disclosure Statement
We declare that there are no relevant financial or non-financial competing interests to disclose.
Supplementary Material for “Sparse Asymptotic PCA: Identifying Sparse Latent Factors Across Time Horizon in High-Dimensional Time Series"
A Examples of Sparse Factors in Time Horizons
  1. (i)
  2. (ii)
    Disaggregate Business Cycle Analysis. Gregory and Head (1999) found that cross-country variations have common components, referred to as global shocks. In addition, each country also experiences country-specific shocks. It is clear that the global and country-specific shocks are reflected by the common factors ${\mathbf{f}}_{t}$’s and the idiosyncratic terms ${\mathbf{e}}_{t}$’s in Model (2.2), respectively. Under the proposed framework, sparse factors over a certain time period correspond to situations where global shocks do not play a dominant role in driving the cyclical variations in a country’s economy. Conversely, nonzero factors will inspire researchers to discover the possible reasons for significant global shocks during the corresponding time period.
  3. (iii)
  4. (iv)
    Monitoring and Forecasting. Stock and Watson (1998) and Forni et al. (2000b) have demonstrated that factor models provide an effective framework for monitoring economic activities, as large-dimensional economic panel data can often be summarized by a small number of factors. In this context, business cycles can be modeled through the co-movements of economic variables, which correspond to the common factors ${\mathbf{f}}_{t}$ in Model (2.2) of the main text. When these factors exhibit sparsity, our approach allows us to track economic activities by identifying systematic risks over time. Furthermore, periods of heightened volatility in the economic panel may produce nonzero factors that are associated with specific events that influence the overall system. Additionally, sparse common factors (diffusion indices) extracted from a large panel of macroeconomic variables may also improve forecasting accuracy when combined with methods such as those proposed by Gao and Tsay (2024).
B Some Tables and Figures for Numerical Studies
In this section, we provide additional tables and figures that support the numerical analysis in the main article. Figure S.1 presents the time plots of the estimated risk factors over the time horizon of the dataset analyzed in Section 5, obtained using PCA. Figure S.2 plots the the empirical histograms of the estimated loading of $\sqrt{T}(\widehat{\boldsymbol{\lambda}}_{1}-{\mathbf{H}}_{s}\boldsymbol{\lambda}_{1})$ for the example of one-factor case, where we set $(N,T)=(300,500)$ and $s=T/10$. Table S.I reports the estimation accuracy of factors when the number of factors $r=3$ for the example of the multi-factor case. Table S.II presents the empirical recovery rate of the sparsity in the factor processes when the number of factors $r=3$ for the multi-factor case. Table S.III reports the empirical probabilities (EP) of determining the number of factors using the ratio-based method in (2.12) of the main article when the number of factors $r=3$. Table S.IV presents the empirical probabilities (EP) of determining the sparsity parameter using the information criterion in (3.9) when the number of factors $r=1$ to verify the accuracy of the proposed criterion.
In Table S.V, we outline the categorization of 332 Stocks in the S&P 500 Index by Industry Sector, which are used in the empirical analysis. Table S.VI presents the nine significant systematic risks over the time horizon from January 1 2004 to December 31, 2016, identified by the proposed sparse factor modeling approach.
Figure S.1: Time plots of the estimated risk factors across the time horizon using the Principal Component Analysis method. The data consists of the daily stock returns for the balanced panel of S&P 500 stocks from January 1st 2004 to December 31st 2016.
Figure S.2: Histograms of the estimated loading of $\sqrt{T}(\widehat{\boldsymbol{\lambda}}_{1}-{\mathbf{H}}_{s}\boldsymbol{\lambda}_{1})$, where we set $(N,T)=(300,500)$ and $s=\left\lceil\sqrt{T}\right\rceil$. The settings of the parameters are the same as those in Table 1 of the main text. The curve represents the density of normal distribution where the standard errors are estimated by the sample versions using the residuals. 2000 replications are used in the experiments.